OpenStudy (anonymous):

what is the equation of the tangent to the curve 3x^2 +3y^2 +2xy -2x -2y=0 at the point (0,0)?

OpenStudy (anonymous):

use the chain rule to get the derivative: \[\frac{d}{dx}(3x^2+3y^2+2xy-2x-2y) = 0\]\[6x+6y\frac{dy}{dx}+2y+2x\frac{dy}{dx}-2-2\frac{dy}{dx} = 0\] Solving for dy/dx gives: \[\frac{dy}{dx} = -\frac{6x+2y-2}{6y+2x-2}\] plugging in (0,0) gives a slope of -1. then use point slope form to get the line.

OpenStudy (anonymous):

thanks!! you are PRO!!